OpenStudy (sabrina1):

Find an equation for the line passing through the point (6, 1) and perpendicular to the line y=2x-14

OpenStudy (anonymous):

1 = -.5(6) + b 1 = -3 + b 4 = b y = -1/2x + 4

OpenStudy (anonymous):

use pt slope rule and use negative reciprocal y-y1=m(x-x1) that will give you the equation of your line and to get a perpendicular slope just use the negative reciprocal

OpenStudy (anonymous):

so y-1=-1/2(x-6) y-1=-1/2x+3 y=-1/2x+4